Three forces (2i^+3j^-2k^)N, (-5i^-5j^+2k^)N and (6i^+5j^-3k)N are working simultaneously on body which moves from the position (in meter) (2,5,-4) to the position (in meter) (-4,5,-3). Calculate the amount of work done.
1
Expert's answer
2020-08-31T12:47:52-0400

The net force acting on a body

"{\bf F}={\bf F}_1+{\bf F}_2+{\bf F}_3"

"{\bf F}=2\hat {\bf i}+3\hat {\bf j}-2\hat {\bf k}-5\hat {\bf i}-5\hat {\bf j}+2\hat {\bf k}+6\hat {\bf i}+5\hat {\bf j}-3\hat {\bf k}\\
=3\hat {\bf i}+3\hat {\bf j}-3\hat {\bf k}"

The displacement of the body

"{\bf d}={\bf r}_2-{\bf r}_1\\
=(-4\hat {\bf i}+5\hat {\bf j}-3\hat {\bf k})-(2\hat {\bf i}+5\hat {\bf j}-4\hat {\bf k})\\
=-6\hat {\bf i}+0\hat {\bf j}+\hat {\bf k}"

The work done

"W={\bf F}\cdot{\bf d}\\
=3\times (-6)+3\times 0+(-3)\times 1=-21 \:\rm J"